A triangle has a perimeter of 161 miles. Each of the two smaller sides of the triangle is two-thirds the length of the longest side. Find the length of each side of the triangle.: A triangle has a perimeter of 161 miles. Each of the two smaller sides of the triangle is two-thirds the length of the longest side. Find the length of each side of the triangle.
Answer by checkley77(3848)

2x/3+2x/3+x=161
4x/3+x=161
(4x+3x)/3=161
7x=3*161
7x=483
x=483/7
x=69 miles for thr longest side.
2/3*69=138/3=46 for the 2 short sides.
Proof:
46+46+69=161
161=161